# On The Fragility of Benchmark Contamination Detection in Reasoning Models

### v1

@@ E.4 Contamination Inflation Comparison
E.5 Log-prob distribution after extensive SFT contamination on LRMs
We provide the log prob distributions for members and non-members of deepseek distill models before and after contamination on Minvera Math and GPQA-Diamond in Fig. 8 and 9. Even though non-members have not been exposed to LRMs during the contamination, the log-prob would also increase, demonstrating that LRMs start to generalize after contamination.
@@ E.5 Log-prob distribution after extensive SFT contamination on LRMs
E.6 Token Embedding Visualization of member and non-member of benchmarks
We provide visualization of token embeddings of member and non member of benchmarks. We can observe that we cannot distinguish between member and non-member using embeddings, which infer that embedding based method is hard to identify the membership. Specifically, we extract the last token from the last layer of the hidden space.
(a) Result on Minerva Math before SFT Contamination
(b) Result on Minerava Math after SFT Contamination
Figure 10: Last token embedding visualization of the final hidden layer before and after SFT contamination. In both cases, member and non-member embeddings are highly overlapped and therefore not distinguishable.
(a) Result on GPQA Diamond before SFT Contamination
(b) Result on GPQA Diamond after SFT Contamination
Figure 11: Last token embedding visualization of the final hidden layer before and after SFT contamination. In both cases, member and non-member embeddings are highly overlapped and therefore not distinguishable.
(a) Result on OlympiadBench before SFT Contamination
(b) Result on OlympiadBench after SFT Contamination
Figure 12: Last token embedding visualization of the final hidden layer before and after SFT contamination. In both cases, member and non-member embeddings are highly overlapped and therefore not distinguishable.
E.7 Additional Results

### v2

@@ E.9 Contamination Inflation Comparison (Stage I: pre-LRM)
E.10 Log-prob distribution after extensive SFT contamination on LRMs (Stage II: post-LRM)
We provide the log prob distributions for members and non-members of deepseek distill models before and after contamination on Minvera Math and GPQA-Diamond in Fig. 13 and 14. Even though non-members have not been exposed to LRMs during the contamination, the log-prob would also increase, demonstrating that LRMs start to generalize after contamination.
@@ E.10 Log-prob distribution after extensive SFT contamination on LRMs (Stage II: post-LRM)
E.11 Applicability in non-math domains (Stage I and II)
We further evaluate our claims on datasets beyond math and science domains for both the pre-LRM and the post-LRM stage. We choose LiveCodeBench v2 (Coding) and MMLU-Pro (General QA). For MMLU-Pro, we only select the ‘health’, ‘history’, ‘law’, ‘other’, ‘philosophy’, and ‘psychology’ domains that are not related to math and science.
We first provide results for the pre-LRM stage (i.e., contamination happens when the base model evolves into LRMs) with Qwen2.5-7B-Instruct as the base model in Table 21. Despite differences in AUROC degradation speed across domains, the trend remains consistent: RL conceals the SFT contamination. Overall, these results indicate that the effect of RL concealments is not limited to the reasoning domains but extends to more general domains as well.
We also provide the AUROC of contamination detection approaches evaluated on advanced LRM that is contaminated in the post-LRM stage (i.e., contamination with CoT on advanced LRMs) in Table 22. The results show that contaminated LRMs leave minimal contamination evidence in coding as well, indicating that the detection performance is near random guess for those reasoning-related domains (e.g., coding, math, science, etc). For domains that require less CoT (e.g., general QA), despite that the AUROC could reach 75% when using the Loss as the detection for the contaminated Deepseek distilled Qwen-14B, the detection performance is much worse compared to results in the pre-LRM stage. Specifically, the AUROC for the Loss detector on MMLU-Pro can reach 93.72%, and LiRA could achieve 79.85% on LiveCodeBench when SFT contamination is applied to the base model, shown in Table 18 in the paper.
Table 21: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO in pre-LRM stage. Results demonstrate that after GRPO, AUROC decreases across both general QA and coding domains and different detection approaches. $\Delta$ measures the difference with the SFT-contaminated model w/o RL. Higher AUROC, better detection performance. The base model here is Qwen2.5-7B-Instruct.
Contamination Detection Methods Training Stages MMLU-Pro (Non-STEM domains) LiveCodeBench Avg. $\Delta$
Reference based
LiRA (Mireshghallah et al., 2022) Before RL 96.68 79.85 88.26 +0.00
RL w/ clean 93.85 71.12 82.48 -5.78
Ref (Carlini et al., 2021) Before RL 86.25 69.51 77.88 +0.00
RL w/ clean 81.67 66.68 74.17 -3.71
Reference free
Loss (Carlini et al., 2021) Before RL 93.72 64.87 79.29 +0.00
RL w/ clean 88.69 54.26 71.47 -7.82
Min–K (Shi et al., 2023) Before RL 93.43 64.12 78.77 +0.00
RL w/ clean 88.54 54.29 71.41 -7.36
Max–K (Maini et al., 2024) Before RL 82.27 63.75 73.01 +0.00
RL w/ clean 59.48 49.80 54.64 -18.37
Table 22: AUROC (%) of contamination detection approaches evaluated on advanced LRMs contaminated with CoT in both general QA and coding domain benchmarks in post-LRM stage.
Contamination Detection Methods Init Models MMLU-Pro (Non-STEM domains) LiveCodeBench Avg.
Reference based
LiRA (Mireshghallah et al., 2022) DS Llama-8B 72.85 61.37 67.11
DS Qwen-14B 74.75 58.78 66.77
Ref (Carlini et al., 2021) DS Llama-8B 60.77 58.75 59.76
DS Qwen-14B 63.43 58.17 60.80
Reference free
Loss (Carlini et al., 2021) DS Llama-8B 69.56 57.42 63.49
DS Qwen-14B 75.83 56.52 66.18
Min–K (Shi et al., 2023) DS Llama-8B 68.12 56.59 62.36
DS Qwen-14B 76.09 56.08 66.09
Max–K (Maini et al., 2024) DS Llama-8B 61.07 56.33 58.70
DS Qwen-14B 66.88 54.80 60.84
@@ Appendix F Computation resources
Appendix G Use of the LLM
We only use LLM to aid paper writing and retrieve related works.